# Patch-Consistent Optical Translation Across Sensors: Large-Scale Denoising Diffusion with Heterogeneous Change Detection as a Use Case

### v2

@@ Abstract
In the field of remote sensing, the challenge of comparing images captured by disparate sensors is a common obstacle. This requires image translation—converting imagery from one sensor domain to another while preserving the original content. Denoising Diffusion Implicit Models (DDIM) are potential state-of-the-art solutions for such domain translation due to their proven superiority in multiple image-to-image translation tasks in classic computer vision. However, these models struggle with large-scale multi-patch imagery, often focusing solely on small patches and resulting in inconsistencies across the full image. To overcome these limitations, we propose a novel method that leverages DDIM for effective optical image translation over large areas. Our approach is tailored to super-resolve large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. Extensive experiments with a dataset of paired Sentinel-II and Planet Dove images show that our approach provides precise domain adaptation and artifact reduction. Our technique preserves the image content while also improving radiometric (color) accuracy and feature representations. The outcome is a high-resolution large-scale image with consistent patches, vital for applications such as heterogeneous change detection (HCD). We present a unique training and testing algorithm rooted in DDIMs, a thorough image quality assessment, and a comparative study against the standard classifier-free guided DDIM framework and five other leading methods. The efficacy of our approach is further demonstrated by substantial enhancements in HCD tasks performed in the urban settings of Beirut, Lebanon, and Austin, USA.
@@ I Introduction
Image-to-image translation (I2I) is the task of transforming an image belonging to one data domain to another domain while preserving at least partially its original content. This is useful for various applications such as super-resolution, inpainting, image decompression, and domain adaptation [Saharia2021, Dhariwal2021, I2ISeo2023, Park2019, I2IIsmael2023, I2ITasar2020, I2IZhang2023, I2ISokolov2023]. In the context of remote sensing, I2I is useful when images acquired from different sensors need to be compared. For example, a deep learning model that learns from paired images can translate a low-resolution domain $A$ optical image to a high-resolution domain $B$ optical image or vice versa. Thus, synthetic versions of high-resolution images from low-resolution ones can be created, which are useful when the original high-resolution images are not available, too costly to acquire, or plagued by unfavorable weather conditions. However, I2I in this context is challenging because of the differences in spatial and spectral characteristics, noise levels, and geometric distortions that are particular to each optical sensor. Therefore, an I2I model needs to learn how to preserve the content and structure of the source image while adapting to the radiometric properties and spatial resolution of the target domain.
We address the limitations of existing Image-to-Image (I2I) methods [CDFang2023, I2IZhang2023, I2ISeo2023, CDLi2021, Xiao2023] in remote sensing imaging, which are primarily designed for patch-level translation and fail to consider the variability of optical images influenced by various factors. Notably, our work is the first to propose an algorithm that integrates large-scale domain adaptation and super-resolution into a single task using Denoising Diffusion Models (DDM) [Ho2020].
Our novel optical-to-optical translation approach leverages the strengths of DDMs. These models have demonstrated exceptional image generation capabilities, surpassing adversarial models without encountering their common training issues, such as mode-collapse [mansourifar_i3eaccess2022]. The method hereby proposed exclusively employs Denoising Diffusion Implicit Models (DDIM), which are advantageous due to their reduced required number of model evaluations when compared to Denoising Diffusion Probabilistic Models (DDPM) and due to their deterministic behavior in intermediate inference steps. This approach ensures efficient and consistent domain-robust large-scale I2I translation, setting our method apart from existing techniques.
The proposed method differs from the standard classifier-free guided DDIM framework commonly used in traditional computer vision tasks [ho2022classifierfree, Saharia2021, Dhariwal2021, Song2020, Nichol2021, Wang2022], as, during inference, it seems to get confused by the high variability of colors between training patches even when their features are very similar. This causes the generation of patches with whose overall colorization (radiometry) is highly inaccurate. Our method tackles this problem by using color-standardization pre- and post-processing procedures carefully tailored to the optical-to-optical translation task.
Moreover, due to its stochastic nature, denoising diffusion models produce different outputs for a given input [Ho2020]. Even when not adding noise in between reverse diffusion steps, as in DDIM [Song2020], different choices of the initial random noise matrix result in different outputs for the same input. This characteristic, while crucial for tasks like inpainting and unconditional generation, can bring uncertainty about the consistency of generations in the context of I2I for remote sensing. In our experiments, we observed that the randomness brought by the initial noise matrix translates into the creation of a few but non-negligible low-quality generations caused by sub-optimal choices of initial noise conditions, leading to inconsistent feature generations across neighboring patches. To alleviate this, we propose a modification in the DDIM inference procedure that optimizes the choice of the initial random noise matrix.
@@
As experiments, we apply our DDM-based method to generate synthetic high-resolution images for eight regions of interest: Beirut, Austin, Tlaquepaque, and five other locations. Our I2I method greatly improves the image quality when using Sentinel-II as input, compared to the original Sentinel-II images and to synthetic images generated by other state-of-the-art models. To test the method in a practical application, we perform two heterogeneous change detection (HCD) experiments. HCD is a technique that compares images acquired from different sensors or different modes of image acquisition to identify changes in a specific area. Valuable for quick disaster damage assessment, it is, however, challenging to achieve accurate and consistent results due to the differences in image characteristics and quality [Chini2023, Touati2020, Zhang2022, WangCD2022, Lv2022, Wu2022]. The presented HCD tests show how our proposed I2I method can alleviate the effect of such differences. These experiments are performed on two regions: Beirut, Lebanon, and Austin, Texas, using a targetless change detection algorithm based on classical image processing tools. We compare HCD results obtained with and without synthetic images. The results demonstrate that the synthetic images generated by the proposed I2I method improve the HCD performance by greatly reducing false alarms and enhancing the change regions. Moreover, the synthetic images generated for this test are realistic, detailed, consistent, and with better contrast and clarity compared to the original low-resolution images.
@@
    We popose original training and testing procedures based on denoising diffusion models that are able to translate optical images from a low spatial resolution domain to a higher one over large areas while maintaining high inter-patch and input-output consistency.
@@
    The presentation of image quality comparisons between the proposed method and other cutting-edge solutions, together with ablation experiments, over hundreds of square kilometers of imagery, that highlights the contribution of our approach in contrast to the classical classifier-free guided DDIM one [ho2022classifierfree].
@@
    The presentation of heterogeneous change detection tests in Beirut, Lebanon, and Austin, USA, indicating the contribution of our proposed I2I translation method to heterogeneous change detection performance.
@@
Figure 1: Surface reflectance imagery from Beirut, Lebanon, of low and high spatial resolutions, captured at similar time intervals, from Sentinel-II (10m) and Planet Dove (3m) sensors. Figures 1a and 1c are the pre and post-event images from Sentinel-II, respectively. Figures 1b and 1d are the pre and post-event images from Planet Dove, respectively.
II Proposed Method
Our method tackles the challenging task of large-scale patch-consistent optical-to-optical image translation from a lower spatial resolution sensor domain to a higher spatial resolution one. While spatial resolution is the most noticeable difference between such domains, there are also other disparities [CuiColor2020] such as radiometric fidelity and consistency, illumination conditions, and other factors such as sensor spectral sensitivity that increases furthermore the radiometric discrepancy between the two domains. The method hereby described leverages denoising probabilistic diffusion models (DDPMs) to not only virtually increase the spatial resolution but also to align subtle yet significant differences between the two domains. This requires a training dataset with pairs of co-registered low spatial resolution (LR) domain $A$ and high spatial resolution (HR) domain $B$ images, showing the same scenes taken with different sensors, to ensure proper domain adaptation.
The full method generates large multi-patch images with highly consistent and color-accurate patches, showing significant visual resolution improvement compared to lower-resolution domain images while preserving the original content (see Section IV). The model architecture and the associated hyperparameters used for our experiments are discussed in Appendix B. Details about the denoising diffusion working principles are presented in Appendix C.
@@ II-A Training
Our training procedure diverges from the conventional DDPM by incorporating classifier-free guidance training [ho2022classifierfree]. This addition enforces alignment between input and output patches. Additionally, we employ pre and post-processing techniques: Whitening and Coloring, respectively. These are critical as they prevent the diffusion network from being confounded by the high variability in patch coloration found within extensive training datasets, particularly when patches have highly similar content. Without these measures, the network fails to accurately predict the channel mean values for the output patches, as shown in Section IV. Details of the implementation of DDPM, DDIM, and Classifier-Free Guidance are respectively presented in Appendices C-A, C-B, and C-C.
Algorithm 1 defines the training step for a single pair of low spatial resolution patch from domain $A$, $\bm{I}_{\text{in}}^{\text{LR}}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and high resolution patch, $\bm{I}_{\text{out}}^{\text{HR}}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, from domain $B$. The goal is to train a model $\bm{\theta}$ that is able to convert $\bm{I}_{\text{in}}^{\text{LR, local}}$ into $\bm{I}_{\text{out}}^{\text{HR}}$. The training step is illustrated by Figure 2. The input and output images have the same pixel resolution but differ in spatial resolution, i.e., $\bm{I}_{\text{in}}^{\text{LR, local}}$ is up-scaled with cubic interpolation to have the pixel resolution of $\bm{I}_{\text{out}}^{\text{HR}}$, while covering the exact same area and the similar features. Thus, patches aligned in time are required for training to work.
Line 1 of the algorithm defines $\bm{I}_{\text{in}}^{\text{LR}}$ as the channel-wise stacking of $\bm{I}_{\text{in}}^{\text{LR,local}}$ and $\bm{I}_{\text{in}}^{\text{LR, global}}$. The matrix $\bm{I}_{\text{in}}^{\text{LR,local}}$ is defined as the region of interest (ROI) of the low-resolution image from domain $A$, which is the patch to be translated to domain $B$ in inference. $\bm{I}_{\text{in}}^{\text{LR, global}}$ represents the surroundings of the region of interest, down-sampled to have the same dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$. $\bm{I}_{\text{in}}^{\text{LR, global}}$ is defined to contain four times the visible area of $\bm{I}_{\text{in}}^{\text{LR,local}}$, i.e., the area that is contained in $\bm{I}_{\text{in}}^{\text{LR,local}}$ plus three additional equally sized regions in its vicinity, thus depicting north, east and southeast patches that are connected to $\bm{I}_{\text{in}}^{\text{LR,local}}$. This is done to give more context to the model of what exactly lies around the input patch.
@@
Line 2 samples a random variable $u$ from a Bernoulli distribution of probability $p_{\text{uncond}}$. If it turns out to be $1$, the model behaves as an unconditional diffusion model for the current training step (Line 4), as in [ho2022classifierfree]. Otherwise, it behaves as a conditional diffusion model. ${\O}$ is the null-condition matrix (like in [ho2022classifierfree]). To ensure that ${\O}$ is correctly interpreted by the model, we set it as a matrix whose all elements are equal to $-2$, which is outside the range of values of input pixels $[-1,1]$. Thus, the model does not interpret the null condition as a regular input image and vice-versa. Lines 6 and 8 define the conditioning input from domain $A$, $\bm{x}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and the domain $B$ ground truth $\bm{y}_{0}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, respectively. They are obtained by applying the function $\mathbf{W}$, defined in Algorithm 2, to $\bm{I}_{\text{in}}^{\text{LR}}$ and $\bm{I}_{\text{out}}^{\text{HR}}$. This function extracts out the overall color information from an input image, i.e., it zeroes the mean of each channel. A normalization between $[-1,1]$ follows. $\mathbf{W}$ also returns the previous means, minimum, and maximum values extracted from the input, which can be used for reverting the whitening process during inference. However, during training, these values are not used. Lines 9 to 13 of Algorithm 1 are as in the original conditional DDPM training procedure [Saharia2021]. This process is repeated for all existing training patches until convergence.
@@
II-B Inference
The inference procedure, illustrated by Figure 3, aims at maximizing the image generation quality by selecting the best possible initial noise matrix to be fed to the reverse diffusion process. We adapt the classifier-free guided DDIM approach by integrating the same color standardization methods used during training. To further enhance the fidelity of the generated patches to their inputs, we refine the selection process for the initial noise condition input for the reverse diffusion process, ensuring optimal consistency and translation quality. Implementation details for the DDIM inference process are presented in Appendix C-B.
@@ II-B Inference
III Dataset
For the proposed experiments, paired images from satellites Sentinel-II and Planet Dove were gathered. The former is a widely known open-access optical satellite with a spatial resolution of 10 meters, whereas the latter is a class of commercial satellites aimed at producing high-quality images at a spatial resolution of 3 meters. A dataset containing 77 pairs of surface reflectance rasters has been assembled for training. These rasters are parts of or a combination of acquisitions that were captured at similar time periods. They are properly geolocated, georeferenced, and checked for image quality issues. In total, a vast area of 7894 $7894\text{ km}^{2}$ is covered by the training set, which is comparable to the area of Puerto Rico (9104 $9104\text{ km}^{2}$). As the test set for the image-to-image translation algorithm, 8 pairs of images were selected, which, in total, cover an area of 707 $707\text{ km}^{2}$. With these, image quality metrics are extracted for the proposed method and compared to other possible approaches for the I2I translation task.

### v3

# Multi-Sensor Diffusion-Driven Optical Image Translation for Large-Scale Applications

@@ Abstract
Comparing images captured by disparate sensors is a common challenge in remote sensing. This requires image translation—converting imagery from one sensor domain to another while preserving the original content. Denoising Diffusion Implicit Models (DDIM) are potential state-of-the-art solutions for such domain translation due to their proven superiority in multiple image-to-image translation tasks in classic computer vision. However, these models struggle with large-scale multi-patch imagery, often focusing solely on small patches and resulting in inconsistencies across the full image. To overcome these limitations, we propose a novel method that leverages DDIM for effective optical image translation over large areas. Our approach super-resolves large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. Our contributions lie in new forward and reverse diffusion processes, which are crucial for addressing the challenges of large-scale image translation. Extensive experiments with a dataset of paired Sentinel-II and Planet Dove images demonstrate that our approach provides precise domain adaptation and artifact reduction. The proposed method preserves image content while improving radiometric (color) accuracy and feature representation. The outcome is a high-resolution large-scale image with consistent patches, vital for applications such as heterogeneous change detection (HCD). We present a unique training and testing algorithm rooted in DDIMs, a thorough image quality assessment, and a comparative study against the standard classifier-free guided DDIM framework and five other leading methods. The usefulness of our approach is further demonstrated in HCD tasks performed in the urban settings of Beirut, Lebanon, and Austin, USA.
@@ I Introduction
Image-to-image translation (I2I) involves transforming an image from one data domain to another while preserving its original content. This technique is useful for applications such as super-resolution, inpainting, image decompression, and domain adaptation [Saharia2021, Dhariwal2021, I2ISeo2023, Park2019, I2IIsmael2023, I2ITasar2020, I2IZhang2023, I2ISokolov2023].
@@
In remote sensing, I2I is valuable when comparing images from different sensors. For example, a deep learning model can translate a low-resolution domain $A$ optical image to a high-resolution domain $B$ optical image.This creates synthetic high-resolution images from low-resolution ones, which are useful when original high-resolution images are unavailable, costly, or affected by weather conditions. However, I2I in this context is challenging due to differences in spatial and spectral characteristics, noise levels, and geometric distortions specific to each sensor. Therefore, an I2I model must preserve the content and structure of the source image while adapting to the radiometric properties and spatial resolution of the target domain.
Despite advancements in image-to-image (I2I) translation for remote sensing, most literature focuses on generating individual synthetic patches for augmenting training data sets used in downstream tasks [I2ITasar2020, I2IIsmael2023, I2ISokolov2023], or on SAR to Optical translation [CDLi2021, I2IZhang2023, I2ISeo2023]. While highly valuable, these methods do not address the challenge of translating large-scale images across disparate optical sensors with different spatial resolutions, particularly ensuring feature consistency between input and output images. This gap is significant as it hinders the fidelity of the translated images, where preserving salient features without introducing non-existent elements is crucial. Additionally, these methods are usually tested on benchmark datasets or small datasets, which may not adequately demonstrate their effectiveness in real-world scenarios.
An appealing choice for addressing these challenges is the use of Denoising Diffusion Models (DDMs) [Ho2020], which have shown exceptional image generation capabilities in other contexts, surpassing adversarial models without common training issues like mode-collapse [mansourifar_i3eaccess2022]. However, standard diffusion methods seem to struggle with high variability in overall patch colorization across samples with similar features in a vast data set. Also, at times they produce inconsistencies in feature generation across neighboring patches due to their stochastic nature [Ho2020].
To overcome these limitations, we propose a novel method that leverages Denoising Diffusion Implicit Models (DDIM) for effective optical image translation over large areas. Our approach super-resolves large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. We address the issues of high variability in patch colorization and inconsistencies in feature generation by designing novel forward and reverse diffusion procedures. These procedures enhance spatial resolution and align radiometric differences between domains, ensuring high-quality, consistent results.
Notably, this work is the first to integrate large-scale optical domain adaptation and super-resolution into a single task using a DDPM-based method. This integration generates large multi-patch images with consistent and color-accurate patches, showing significant visual resolution improvement while preserving the original content (see Section V). The robust training dataset and scalability of our approach contribute to its state-of-the-art performance, making it suitable for practical applications in remote sensing.
@@
We apply our DDM-based method to generate synthetic high-resolution images for eight regions of interest: Beirut, Austin, Tlaquepaque, and five other locations. In our experiments, Sentinel-II imagery ($10\text{m}^{2}/\text{pixel}$) is translated into synthetic versions of Planet Dove imagery ($3\text{m}^{2}/\text{pixel}$). We perform two heterogeneous change detection (HCD) experiments to test the method in a practical application. HCD compares images from different sensors or modes of acquisition to identify changes in a specific area. Valuable for quick disaster damage assessment, achieving accurate and consistent results is challenging due to differences in image characteristics and quality [Chini2023, Touati2020, Zhang2022, WangCD2022, Lv2022, Wu2022].
While we present heterogeneous change detection (HCD) experiments to showcase one application of our image-to-image (I2I) translation algorithm, it is important to note that these experiments are not intended to demonstrate peak performance in HCD. The simplistic image processing algorithm used in these experiments does not exploit the resolution gains from our diffusion-based image translation model. Instead, the purpose is to illustrate the potential use of our I2I translation method in HCD. Developing a learning-based HCD method that fully leverages the added resolution is beyond the scope of this article.
@@
    We propose original training and testing procedures based on denoising diffusion models that translate optical images from a low spatial resolution domain to a higher one over large areas while maintaining high inter-patch and input-output consistency.
@@
    We present image quality comparisons between our method and other cutting-edge solutions, along with ablation experiments over hundreds of square kilometers of imagery, highlighting the contributions of our approach compared to traditional denoising diffusion.
@@
    We present heterogeneous change detection tests in Beirut, Lebanon, and Austin, USA, indicating the potential usefulness of our method in HCD tasks.
II Related Works
II-1 Denoising Diffusion Models
II-2 Image-to-Image Translation
@@
Figure 1: Surface reflectance imagery from Beirut, Lebanon, of low and high spatial resolutions, captured at similar time intervals, from Sentinel-II (10m) and Planet Dove (3m) sensors. Figures 1a and 1c are the pre- and post-event images from Sentinel-II, respectively. Figures 1b and 1d are the pre- and post-event images from Planet Dove, respectively.
III Proposed Method
Our method addresses the challenging task of large-scale, patch-consistent optical-to-optical image translation from a lower to a higher spatial resolution sensor domain. While spatial resolution is the most noticeable difference, other disparities such as radiometric fidelity, illumination conditions, and sensor spectral sensitivity also increase the radiometric discrepancy between the two domains [CuiColor2020].
To tackle these challenges, we leverage Denoising Probabilistic Diffusion Models (DDPMs) to not only enhance spatial resolution but also align subtle yet significant differences between the two domains, ensuring improved radiometric fidelity and consistency. This requires a training dataset with pairs of co-registered low spatial resolution (LR) domain $A$ and high spatial resolution (HR) domain $B$ images, capturing the same scenes with different sensors for proper domain adaptation.
The method hereby proposed exclusively employs Denoising Diffusion Implicit Models (DDIM) [Song2020], which are advantageous due to their reduced number of model forward passes during inference and their deterministic behavior in intermediate steps. However, our approach differs from the standard classifier-free guided DDIM framework commonly used in traditional computer vision tasks [ho2022classifierfree, Saharia2021, Dhariwal2021, Song2020, Nichol2021, Wang2022]. The standard framework struggles with the high variability in patch coloration within large training sets, leading to inaccurate colorization in generated patches.
Additionally, due to the stochastic nature of denoising diffusion models, different initial noise matrices can result in varying outputs for the same input [Ho2020]. This randomness can cause inconsistencies in feature generation across neighboring patches, which is problematic for image-to-image translation in remote sensing.
The proposed method addresses these deficiencies by implementing novel forward and reverse diffusion procedures, ensuring high-quality, consistent results.
The model architecture and the associated hyperparameters used for our experiments are discussed in Appendix A. Details about the denoising diffusion working principles are presented in Appendix B.
@@ III-A Training
Our training procedure diverges from the conventional DDPM by incorporating modifications to the forward diffusion process through Whitening and Coloring techniques. These modifications are critical as they prevent the diffusion network from being confounded by the high variability in patch coloration found within extensive training datasets, particularly when patches have highly similar content. Without these measures, the network fails to accurately predict the channel mean values for the output patches, as shown in Section V. Details of the implementation of DDPM and DDIM are presented in Appendices B-A and B-B.
Algorithm Description
@@
Algorithm 1 defines the training step for a single pair of low spatial resolution patch from domain $A$, $\bm{I}_{\text{in}}^{\text{LR}}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and high resolution patch, $\bm{I}_{\text{out}}^{\text{HR}}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, from domain $B$. The goal is to train a model $\bm{\theta}$ that is able to convert $\bm{I}_{\text{in}}^{\text{LR, local}}$ into $\bm{I}_{\text{out}}^{\text{HR}}$. The training step is illustrated by Figure 2. The input and output images have the same pixel resolution but differ in spatial resolution, i.e., $\bm{I}_{\text{in}}^{\text{LR, local}}$ is up-scaled with cubic interpolation to have the pixel resolution of $\bm{I}_{\text{out}}^{\text{HR}}$, while covering the exact same area and the similar features. Thus, patches aligned in time are required for training to work. Line 1 of the algorithm defines $\bm{I}_{\text{in}}^{\text{LR}}$ as the channel-wise stacking of $\bm{I}_{\text{in}}^{\text{LR,local}}$ and $\bm{I}_{\text{in}}^{\text{LR, global}}$. The matrix $\bm{I}_{\text{in}}^{\text{LR,local}}$ is defined as the region of interest (ROI) of the low-resolution image from domain $A$, which is the patch to be translated to domain $B$ in inference. $\bm{I}_{\text{in}}^{\text{LR, global}}$ represents the surroundings of the region of interest, down-sampled to have the same dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$. $\bm{I}_{\text{in}}^{\text{LR, global}}$ is defined to contain four times the visible area of $\bm{I}_{\text{in}}^{\text{LR,local}}$, i.e., the area that is contained in $\bm{I}_{\text{in}}^{\text{LR,local}}$ plus three additional equally sized regions in its vicinity, thus depicting north, east and southeast patches that are connected to $\bm{I}_{\text{in}}^{\text{LR,local}}$. This is done to give more context to the model of what exactly lies around the input patch. Line 2 samples a random variable $u$ from a Bernoulli distribution of probability $p_{\text{uncond}}$. If it turns out to be $1$, the model behaves as an unconditional diffusion model for the current training step (Line 4), as in [ho2022classifierfree]. Otherwise, it behaves as a conditional diffusion model. ${\O}$ is the null-condition matrix (like in [ho2022classifierfree]). To ensure that ${\O}$ is correctly interpreted by the model, we set it as a matrix whose all elements are equal to $-2$, which is outside the range of values of input pixels $[-1,1]$. Thus, the model does not interpret the null condition as a regular input image and vice-versa. Lines 6 and 8 define the conditioning input from domain $A$, $\bm{x}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and the domain $B$ ground truth $\bm{y}_{0}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, respectively. They are obtained by applying the function $\mathbf{W}$, defined in Algorithm 2, to $\bm{I}_{\text{in}}^{\text{LR}}$ and $\bm{I}_{\text{out}}^{\text{HR}}$. This function extracts out the overall color information from an input image, i.e., it zeroes the mean of each channel. A normalization between $[-1,1]$ follows. $\mathbf{W}$ also returns the previous means, minimum, and maximum values extracted from the input, which can be used for reverting the whitening process during inference. However, during training, these values are not used. Lines 9 to 13 of Algorithm 1 are as in the original conditional DDPM training procedure [Saharia2021]. This process is repeated for all existing training patches until convergence.
@@
III-B Inference
The inference procedure illustrated by Figure 3 aims to maximize the image generation quality by selecting the best possible initial noise matrix to be fed to the reverse diffusion process. We propose a novel reverse diffusion procedure by integrating the same Whitening and Coloring techniques used during training. To further enhance the fidelity of the generated patches to their inputs, we refine the selection process for the initial noise condition input for the reverse diffusion process, ensuring optimal consistency and translation quality. Implementation details for the DDIM inference process are presented in Appendix B-B.
@@ III-B Inference
IV Dataset
For the proposed experiments, paired images from satellites Sentinel-II and Planet Dove were gathered. The former is a widely known open-access optical satellite with a spatial resolution of $10\text{m}^{2}/\text{pixel}$, whereas the latter is a class of commercial satellites aimed at producing high-quality images at a spatial resolution of $3\text{m}^{2}/\text{pixel}$. A dataset containing 77 pairs of surface reflectance rasters has been assembled for training. These rasters are parts of or a combination of acquisitions that were captured at similar time periods. They are properly geolocated, georeferenced, and checked for image quality issues. In total, a vast area of 7894 $7894\text{ km}^{2}$ is covered by the training set, which is comparable to the area of Puerto Rico (9104 $9104\text{ km}^{2}$). As the test set for the image-to-image translation algorithm, 8 pairs of images were selected, which, in total, cover an area of 707 $707\text{ km}^{2}$. With these, image quality metrics are extracted for the proposed method and compared to other possible approaches for the I2I translation task.